# Learning from the electronic structure of molecules across the periodic table

### v1

@@ 1 Introduction
    Develop HELM - “Hamiltonian-trained Electronic-structure Learning for Molecules", a scalable Hamiltonian matrix prediction network which enables transfer-learning to energy prediction tasks from a unified atomic orbital basis. HELM is state-of-the-art on the most challenging Hamiltonian matrix benchmarks openly available, and is the first ‘universal’ electronic structure prediction model demonstrated on structure sizes and chemical diversity comparable to those treated by current MLIPs.
@@ 3 Methods
(2) The irreps in the diagonal blocks of $\mathbf{H}$, which represent interactions between orbitals on the same atom have additional parity constraints:
@@
Here, $\ell_{1}^{A}$, $\ell_{2}^{A}$ represent the degree of the interacting orbitals within atom $A$, $\delta$ is the Kronecker delta, and the $\ell_{3}^{A}$s represent their tensor product in the coupled basis. To take advantage of constraints, we compute only unique and non-zero node values.
@@ 3.1 OMol closed-shell 58k
Here, we posit an additional two reasons why naive element-wise losses may fail. First, as the number of heavy elements increases, the distribution of magnitudes in the matrix elements of $\mathbf{H}$ increases disproportionately. The parallel problem in energy prediction has typically been mitigated by computing reference values per atomic element, and using them to rescale the training data. Second, MAE losses are performed component-wise, which may bias the loss computed over otherwise rotationally-identical edges (Wang et al., 2024). To mitigate these effects, we first pre-process the node labels to scale and center the $l=0$ components, which stem roughly from element-specific localized core electron states (see Section˜B.3). This follows a similar process applied in other recent work on Hamiltonian prediction for materials (Haldar et al., 2025; Qian et al., 2025), and serves to effectively homogenize the variance of the data across different elements. Finally, we train with a root-MSE+MSE loss over the labels, which is independent of irrep orientation.
@@ 4.1 Hamiltonian matrix benchmarks
Table 2: Comparison of H matrix element prediction error on the MD17 and $\nabla^{2}$DFT molecule datasets, for SchNOrb (Schütt et al., 2019b), PhiSNet (Unke et al., 2021), QHNet (Yu et al., 2023b), and SPHNet (Luo et al., 2025). In parentheses under each dataset are the average number of atoms per molecule. Element-wise matrix errors are computed as $H^{err}=\frac{1}{N}\sum_{i,j}|H_{ij}^{\text{pred}}-H_{ij}^{\text{ref}}|$. The best performing model is indicated in bold, and the second best is underlined. Benchmark numbers for $\nabla^{2}$DFT are taken from Khrabrov et al. (2024). In the last row, we summarize the MAE energy error ($E^{err}$) computed from $H_{ij}^{\text{pred}}$ and $H_{ij}^{\text{ref}}$ using PySCF.
@@ D.3 Energy prediction on different conformers splits - $\nabla^{2}$DFT
In Table˜3, we evaluated energy errors on matching 2k, 5k, and 10k conformer train/test splits in $\nabla^{2}$DFT, assessing generalization to different conformers of molecules present in the training set. Hamiltonian pretraining significantly reduced energy prediction errors. In Table˜9, we extend this analysis to test generalization to unseen molecules by evaluating each model, trained on 2k, 5k, or 10k splits, across all test splits. Hamiltonian pretraining consistently outperforms direct training on unseen molecules, even when a finetuned model trained on the 2k split is tested on the additional 8k molecules in the 10k split (8.62 meV/atom) compared to an energy-direct model trained on the 10k split and tested on 2k (11.0).
@@ D.4 Extended UMAP embedding analysis
In Figure˜13 and Figure˜14, we show an extended version of Figure˜4 for both the $\nabla^{2}$DFT and OMol_CSH_58k datasets. We extract embeddings from 100 molecules of the train-2k split for the former, and 250 molecules for the latter. Clusters from the “direct" model indicate that the ability to distinguish hydrogen and carbons atoms has been learned, consistent with the ubiquity of these elements in the training data. After Hamiltonian pretraining, a greater variety of features for different atomic elements can be distinguished. The remaining similarities occur between pairs such as C-H and O-F, which may have more structurally similar atomic environments within the training data.

### v2

@@ 1 Introduction
    Develop HELM - “Hamiltonian-trained Electronic-structure Learning for Molecules", a scalable Hamiltonian matrix prediction network which enables transfer-learning to energy prediction tasks from a unified atomic orbital basis. HELM is state-of-the-art on the most challenging Hamiltonian matrix benchmarks openly available, and is the first ‘universal’ electronic structure prediction model demonstrated on structure sizes and chemical diversity comparable to those treated by current MLIPs. HELM’s scalability to large system sizes, basis sets, and heavy elements is enabled by a combination of SO(2) convolutions (Passaro & Zitnick, 2023), equivariant data scaling, and parity-expansion layers to enforce diagonal sub-matrix symmetry.
@@ 3 Methods
(2) The irreps in the diagonal blocks of $\mathbf{H}$, which represent interactions between orbitals on the same atom have additional parity constraints (relevant ablations are in Section˜C.3):
@@
Here, $\ell_{1}^{A}$, $\ell_{2}^{A}$ represent the degree of the interacting orbitals within atom $A$, $\delta$ is the Kronecker delta, and the $\ell_{3}^{A}$s represent their tensor product in the coupled basis. To take advantage of these constraints, we compute only unique and non-zero node values.
@@ 3.1 OMol closed-shell 58k
Here, we posit an additional two reasons why naive element-wise losses may fail. First, as the number of heavy elements increases, the distribution of magnitudes in the matrix elements of $\mathbf{H}$ increases disproportionately. The parallel problem in energy prediction has typically been mitigated by computing reference values per atomic element, and using them to rescale the training data. Second, MAE losses are performed component-wise, which may bias the loss computed over otherwise rotationally-identical edges (Wang et al., 2024). To mitigate these effects, we first pre-process the node labels to scale and center the $l=0$ components, which stem roughly from element-specific localized core electron states (see Section˜B.3). This follows a similar process applied in other recent work on Hamiltonian prediction for materials (Haldar et al., 2025; Qian et al., 2025), and serves to effectively homogenize the variance of the data across different elements, resulting in a better correlation between matrix element errors and errors in downstream quantities (an ablation study of this effect is in Section˜C.3). Finally, we train with a root-MSE+MSE loss over the labels, which is independent of irrep orientation.
@@ 4.1 Hamiltonian matrix benchmarks
Table 2: Comparison of H matrix element prediction error on the MD17 and $\nabla^{2}$DFT molecule datasets, for SchNOrb (Schütt et al., 2019b), PhiSNet (Unke et al., 2021), QHNet (Yu et al., 2023b), and SPHNet (Luo et al., 2025). In parentheses under each dataset are the average number of atoms per molecule. Element-wise matrix errors are computed as $H^{err}=\frac{1}{N}\sum_{i,j}|H_{ij}^{\text{pred}}-H_{ij}^{\text{ref}}|$. The best performing model is indicated in bold, and the second best is underlined. Benchmark numbers for $\nabla^{2}$DFT are taken from Khrabrov et al. (2024). In the last row, we summarize the MAE energy error ($E^{err}$) computed from $H_{ij}^{\text{pred}}$ and $H_{ij}^{\text{ref}}$ using PySCF, with the functional specific to each dataset.
@@ C.2 Training parameters
C.3 Ablations on node-parity expansion and $l=0$ scaling
Here we preform ablation studies to understand the impact of two specific features of HELM: (1) the node parity expansion blocks described in Section˜3 (‘Hamiltonian output head’), and (2) the $l=0$ component scaling described in Section˜B.3. In Table˜9 we report the results of these ablations using the $\nabla^{2}$DFT 2k train/test splits. All of the training here was done with a fixed number of epochs (500) and a cosine scheduler. We use a fixed number of epochs specifically to compare between these different settings - this is generally not long enough to converge the training (which was done using a ReduceLRonPlateau scheduler for the results in Table˜2), so we also report the error on the train split.
$l=0$ Scaling Node-Parity Expansion Train Test
$H^{err}$ $E^{err}$ $H^{err}$ $E^{err}$
$\times$ $\times$ 201.21 169,026.08 324.35 329,923.53
$\times$ $\checkmark$ 95.73 1,163.54 101.73 1,530.83
$\checkmark$ $\times$ 115.35 837.07 120.18 973.47
$\checkmark$ $\checkmark$ 92.00 540.19 96.27 706.39
Table 9: Effect of node-parity reduction and $l=0$ referencing scheme on matrix element error, evaluated using the $\nabla^{2}$DFT 2k train/test splits. $H^{err}$ is in units of $\times 10^{-6}E_{h}$, while $E^{err}$ is in units of meV. The best case performance in each column is in bold.
The node parity expansion block only impacts the error within diagonal blocks of the Hamiltonian. These diagonal blocks contain the largest-magnitude elements (resulting from nearsightedness of orbital interactions), and contribute disproportionately towards the eigenvalue error and total energy prediction. Therefore, improving the loss over diagonal orbital blocks has a significant impact on the total energy computed from the matrices, bringing the error down from 329,923 meV to 1,530 meV. The $l=0$ component scaling procedure serves to reduce the numerical range of the data from the order of $\sim$100 to the order of $\sim$1 (see Figure˜7 in Section˜B.3), thus making it easier to learn with a properly-normalized model. This brings the energy error down further to 706 meV. Together, these two architectural choices contribute significantly towards the improved performance of HELM over other models in Table˜2.
@@ D.2 OMol_CSH_58k Hamiltonian predictions
To further analyze the utility of the predicted matrices, we use them to re-compute the electron density with PySCF (see Figure˜13). Visually, the electron density isosurfaces computed for all three datasets are nearly indistinguishable from those computed directly from the reference Hamiltonian matrices.
@@
Figure 13: Isosurfaces of electron density visualized for a random molecule taken from the (a) QM7 (isovalues: 0.1, 0.2, 0.5 $\mathrm{e}^{-}/a_{0}^{3}$), (b) $\nabla^{2}$DFT Test-2k (isovalues: 0.05, 0.2 $\mathrm{e}^{-}/a_{0}^{3}$) and (c) OMol_CSH_58k (isovalues: 0.03 $\mathrm{e}^{-}/a_{0}^{3}$) test splits. In each case, the upper row shows the electron density plotted with the label Hamiltonian as input, and the bottom row shows the one predicted from HELM’s output.
@@ D.3 Energy prediction on different conformers splits - $\nabla^{2}$DFT
In Table˜3, we evaluated energy errors on matching 2k, 5k, and 10k conformer train/test splits in $\nabla^{2}$DFT, assessing generalization to different conformers of molecules present in the training set. Hamiltonian pretraining significantly reduced energy prediction errors. In Table˜10, we extend this analysis to test generalization to unseen molecules by evaluating each model, trained on 2k, 5k, or 10k splits, across all test splits. Hamiltonian pretraining consistently outperforms direct training on unseen molecules, even when a finetuned model trained on the 2k split is tested on the additional 8k molecules in the 10k split (8.62 meV/atom) compared to an energy-direct model trained on the 10k split and tested on 2k (11.0).
@@ D.4 Extended UMAP embedding analysis
In Figure˜14 and Figure˜15, we show an extended version of Figure˜4 for both the $\nabla^{2}$DFT and OMol_CSH_58k datasets. We extract embeddings from 100 molecules of the train-2k split for the former, and 250 molecules for the latter. Clusters from the “direct" model reflect the ubiquity of carbon and hydrogen elements in the training data. After Hamiltonian pretraining, a greater variety of features for different atomic elements can be distinguished. The remaining similarities occur between pairs such as C-H and O-F, which may have more structurally similar atomic environments within the training data.